A researcher computes a 2 x 3 chi-square test for independence. What is the critical value for this test at a.05 level of significance?
Answer #1

Please let me know in case you've doubts. Here' the answe below.  

Lets solve this in 2 ways.

First by using the Chi-Square cumulative distribution table.

1.

The cumulative area is 0.95 ( 1-.05 = .95)

The Chi-Square can be got from the ChiSquare table. Open it look for df = (k1-1)*(k2-1), i.e. (2-1)*(3-1) = 1*3 = 3

cumualtive probability = .95.

So, look for df = 3 ( df1 = 1, df2 = 2), with probability = 0.95

It is Chi-Square is  7.8147

The critical value of 7.8147

2.

The cumulative area is 0.95 ( 1-.05 = .95)

The Chi-Square = CHISQ.INV(0.95, 3)

X^2 = 7.8147

The critical value of 7.8147